OpenStudy (anonymous):

Find the magnitude & direction of the vector JK given points J(-6,-4) & K(-10,-4)

OpenStudy (lgbasallote):

\[\text{magnitude} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\text{direction} = \tan^{1} \left(\frac{y_2 - y_1}{x_2 - x_1}\right)\] does that help?

OpenStudy (lgbasallote):

that's supposed to be \[\huge \text{direction} = \tan^{-1} \left(\frac{y_2 - y_1}{x_2 - x_1} \right)\]

OpenStudy (anonymous):

aha thank you that helps a lot.(:

OpenStudy (lgbasallote):

welcome ^_^